### v2

# GME: Improving Universal Multimodal Retrieval by Multimodal LLMs

Abstract
Universal Multimodal Retrieval (UMR) aims to enable search across various modalities using a unified model, where queries and candidates can consist of pure text, images, or a combination of both. Previous work has attempted to adopt multimodal large language models (MLLMs) to realize UMR using only text data. However, our preliminary experiments demonstrate that more diverse multimodal training data can further unlock the potential of MLLMs. Despite its effectiveness, the existing multimodal training data is highly imbalanced in terms of modality, which motivates us to develop a training data synthesis pipeline and construct a large-scale, high-quality fused-modal training dataset. Based on the synthetic training data, we develop the General Multimodal Embedder (GME), an MLLM-based dense retriever designed for UMR. Furthermore, we construct a comprehensive UMR Benchmark (UMRB) to evaluate the effectiveness of our approach. Experimental results show that our method achieves state-of-the-art performance among existing UMR methods. Last, we provide in-depth analyses of model scaling and training strategies, and perform ablation studies on both the model and synthetic data.
1 Introduction
The growth of multimedia applications necessitates retrieval models that extend beyond traditional text-to-text and text-to-image search [75]. In Universal Multimodal Retrieval (UMR) tasks, both queries and candidates can exist in any modality [39]. Compared to addressing this challenge with separate uni-modal and cross-modal retrievers in a divide-and-conquer pipeline [4], a unified retriever is a more viable option in terms of usability and scalability. Using the dense retrieval paradigm (also known as embedding-based retrieval) [25], a unified model can be trained to project inputs from various modalities into a shared embedding space [75, 74, 22]. In this space, similarity scores are computed between the embeddings of queries and the retrieval collection, facilitating the efficient ranking of the top- candidates. To achieve this, some previous studies have primarily focused on two approaches: (1) designing feature fusion mechanisms for cross-modal retrievers based on the CLIP architecture [39, 66], and (2) incorporating visual plugin modules into optimized text embedding models to achieve unified multimodal representations [75, 74].
Recently, researchers have turned to exploring Multimodal Large Language Models (MLLMs) [35, 65] in UMR. For example, it is shown that training MLLMs with text data alone can generate universal multimodal embeddings with respectable retrieval performance [22]. However, modality-limited training may fail to fully demonstrate the potential of MLLMs in UMR. We believe that incorporating multimodal data composition (as shown in Figure 1) could further enhance the model performance and generalization. Moreover, visual documents (i.e. document screenshots) are increasingly important in UMR tasks, as they not only simplify the pipelines of diverse Retrieval-Augmented Generation (RAG) applications, but also mitigate information loss during modality conversion [41, 12]. However, current UMR models primarily target natural images, neglecting support for this scenario (Table 1).
| Methods | Modeling | Retrieval Setting | |||
| Approach | Training | S&C | Fused | VD | |
| UniVL-DR [39] | CLIP Feat. Fusion | Cross-modal | ✓ | ✗ | ✗ |
| UniIR [66] | CLIP Score Fusion | Multimodal | ✓ | ✓ | ✗ |
| BLIP Feat. Fusion | |||||
| MARVEL [75] | Text Enc.+Plugin | Cross-modal | ✓ | ✗ | ✗ |
| VISTA [74] | Text Enc.+Plugin | Multimodal | ✓ | ✓ | ✗ |
| E5-V [22] | MLLM | Text-only | ✓ | ✓ | ✗ |
| GME (Ours) | MLLM | Multimodal | ✓ | ✓ | ✓ |
To address the aforementioned challenges, we propose the General Multimodal Embedder (GME), an instruction-based embedding framework utilizing MLLMs as the backbone. GME enables retrieval across various modalities in the unified paradigm, including text, images, visual documents, and fused-modal111 We use fuse-modal instead of multimodal to denote the data that contains both text and image to disambiguate from the UMR task. (i.e. image-text composed) contents. Our framework is underpinned by two key techniques: (1) A strategically optimized training data composition for UMR. We categorize UMR tasks into three types: single-modal, cross-modal, and fused-modal (Figure 1). Through extensive experimentation, we analyze how different compositions affect performance (Figure 3) and demonstrate that a balanced mixture of all types yields optimal results. (2) An efficient fused-modal data synthesis pipeline. Recognizing the under-representation of fused-modal data and its potential impact on training effectiveness, we develop a streamlined data synthesis pipeline (§4.2). This approach has successfully generated a comprehensive dataset of 1.1M fused-modal pairs, significantly enhancing our training and model capabilities.
To evaluate the effectiveness of our framework, we compile a comprehensive UMR Benchmark, namely UMRB. This benchmark encompasses tasks from widely recognized retrieval benchmarks in text [55], multimodal [66], and visual document retrieval [12], as well as our newly processed fused-modal retrieval data. We build our models on top of the strong Qwen2-VL series MLLMs [65] and train them on our constructed dataset. Experimental results demonstrate that our model achieves state-of-the-art performance on UMRB. Additionally, we perform in-depth analyses on model scaling, training strategies, and ablation of our synthetic data. Our key contributions are:
We explore strategies to adapt MLLMs into UMR models, and present GME, a powerful embedding model capable of retrieving candidates across different modalities. GME is the first UMR model to deliver visual document retrieval performance on par with specialized models.
We propose a novel data synthesis pipeline for constructing large-scale, fused-modal training data to encounter the scarcity of such training data. This pipeline is more efficient than previous approaches and can be easily extended to other domains.
We compile the UMR benchmark, UMRB, to evaluate a broader range of retrieval tasks compared to existing benchmarks. UMRB categorizes tasks into three types: single-modal, cross-modal, and fused-modal, and offers a comprehensive performance evaluation across them.
2 Related Work
Multimodal Large Language Models
The emergence of Large Language Models (LLMs) has driven significant progress in natural language processing [3, 49], leading to the development of Multimodal LLMs that extend these capabilities to handle multimodal information. Prominent MLLMs such as GPT-4V [48], LLaVa [35, 36], Qwen-VL [65], InternVL [7] and MiniCPM-V [71] have shown promising advancements in multimodal information understanding and reasoning. Typically, an MLLM consists of an LLM, a vision encoder, and a projector that bridges the two components by transforming raw multimodal inputs into vectors compatible with the LLM [72].
Multimodal Retrieval
Early multimodal retrieval tasks focused on single-modal [73] or cross-modal retrieval [61]. Recently, the expansion of multimedia applications and multimodal retrieval-augmented generation (RAG) by MLLMs has created a need for unified multimodal retrieval models for complex scenarios. Existing approaches largely utilize pre-trained models such as CLIP [51] or BLIP [29] for multimodal embedding. For instance, UniVL-DR [39] and UniIR [66] initially encode images and texts separately using CLIP or BLIP encoders, followed by fusion strategies like score fusion to integrate features from both modalities. Additionally, VISTA [74] and MARVEL [75] employ pre-trained text embedding models enhanced with visual plugins to encode composite image-text candidates. However, these methods are typically designed for specific tasks like multimodal document retrieval and lack flexibility to handle diverse multimodal retrieval tasks.
Concurrent with our work, E5-V [22] and VLM2VEC [23] propose fine-tuning MLLMs on single-text (NLI [14]) or vision-centric relevance data, demonstrating their transferability to multimodal retrieval. In this paper, we are the first to explore the fine-tuning of an MLLM-based universal multimodal retriever that can address both visual retrieval tasks and maintain strong text-to-text retrieval capabilities. Moreover, we are the first to extend a unified retrieval model to handle not only natural image retrieval but also text-rich image retrieval [12].
Embedding Models with Pre-trained Language Models
With the advancement of pre-trained Language Models, research in both pure text and Vision-Language Models has focused on building representation models based on these pre-trained language models. In the text retrieval domain, state-of-the-art text embedding models such as Contriver [21], E5 [62], GTE [31], and BGE [68] are all built upon pre-trained language models and have demonstrated impressive generalization and robust performance in text retrieval tasks. Recently, leveraging LLMs combined with supervised fine-tuning (SFT), researchers have developed unified text representation models that fully utilize the text understanding capabilities of LLMs, resulting in models with enhanced performance and generalization [63, 31, 28]. These models typically process user text inputs through LLMs, using the hidden states from the final transformer layer—either through pooling or by selecting the last token—as the final representation. Inspired by the success of universal text embedding models based on text LLMs, researchers have begun to explore the construction of unified multimodal retrieval models using MLLMs [22, 23]. In this paper, we aim to demonstrate through systematic experiments that constructing a truly universal multimodal retrieval model using MLLMs is feasible.
3 Universal Multimodal Retrieval
Current UMR sub-tasks can be categorized into three types based on the modalities of the query and the candidate:
Single-Modal Retrieval: Both the query and the candidate belong to the same modality, such as text-to-text (TT) or image-to-image (II) retrieval scenarios.
Cross-Modal Retrieval: The query and the candidate belong to different modalities, typically text-to-image (TI) retrieval. Unlike most prior work that focuses on natural-style image retrieval, we also consider the retrieval of rich-text images (e.g., images converted from scholarly PDFs). We denote this scenario as text-to-visual document (TVD) retrieval.
Fused-Modal Retrieval: More complicated retrieval tasks involve mixed modalities in queries, candidates, or both. For example, in EVQA [46], both queries and candidates combine text and images.
The visualization of these settings refers to Figure 1.
| Class | Task | Datasets | ||||||
| Single- Modal (17) | TT (16) |
| II (1) | Nights[13] | |||||||
| Cross- Modal (18) | TI (4) |
| TVD (10) |
| IT (4) |
| Fused- Modal (12) | TIT (2) | WebQA[4] EDIS[37] | ||||||
| ITT (5) |
| ITI (2) | FashionIQ[67] CIRR[38] | |||||||
| ITIT (3) | OVEN[20] EVQA[46] INFOSEEK[6] |
3.1 Universal Multimodal Retrieval Benchmark
Based on the aforementioned classification principles, we introduce a new benchmark to comprehensively assess the performance of UMR models. This benchmark comprises 47 evaluation datasets that cover a broad spectrum of multimodal retrieval tasks, and we name it the Universal Multimodal Retrieval Benchmark (UMRB). These evaluation datasets primarily originate from previously constructed datasets tailored for each sub-scenario or sub-task. Specifically, UMRB includes: (1) The BEIR [55] benchmark for text-to-text retrieval scenarios; (2) The M-BEIR [66] dataset for vision-centric retrieval scenarios; (3) Additional fused-modal datasets that not cover by M-BEIR; and (4) text-to-visual document search datasets, such as ViDoRe [12], to extend the coverage of our benchmark and ensure a comprehensive evaluation of model universality. A detailed list of the UMRB datasets is presented in Table 2.
Given the extensive size of UMRB, to expedite our experimental validation and analysis, we have sampled a subset of datasets from each category, constituting a smaller dataset named UMRB-Partial. This subset retains of the total datasets while maintaining evaluation richness. More detailed statistical information about UMRB-Partial can be found in Appendix Table 6.
4 Method
In this section, we present the training framework for developing the General Multimodal Embedder (GME) model. We describe the contrastive learning approach used to train the embedding model. Building on this, we conduct detailed experiments to determine the optimal balance of training data type. Specifically, our experiments demonstrate that diverse data type mixtures significantly enhances the model’s ability to perform retrieval across various modalities. Lastly, recognizing the scarcity of high-quality fused-modal training data, we propose a novel method for automatically synthesizing large-scale, high-quality training data using MLLM.
4.1 GME: General Multimodal Embedder
Model Architecture
We employ a MLLM as the foundation for GME. This model can accept images, text, or image-text pairs as input. Inspired by previous research on text embedding [31, 63], we use the final hidden state of the last token as the representation (or embedding) for the input. Although pre-trained MLLMs possess strong multimodal understanding capabilities, their original training objectives are not optimized for representation learning. Therefore, task-specific fine-tuning (or alignment) is necessary to enhance the model’s representational capacity. Contrastive learning has been shown to effectively train LLMs and MLLMs to produce retrieval embeddings [31, 22].
Contrastive Learning
In our contrastive learning setup, each training instance comprises a query , a relevant candidate , and a set of irrelevant candidates . Both and can be text, images, or image-text pairs, allowing the model to handle diverse data modalities. To tailor the model to various downstream retrieval tasks, we incorporate an instruction tuning method by including a tailored instructional text with each retrieval task. For example, for the Visual Question Answering (VQA) task, the instruction could be: “Retrieve a passage that provides an answer to the given query about the image” guiding the model on how to process and interpret the query for specific objectives.
During training, we input and instruction into the model to obtain the query representation . Similarly, each candidate is input into the model to obtain its representation . The training objective minimizes the cosine distance between and for relevant pairs while maximizing the distance between and for irrelevant pairs. Cosine similarity is employed to measure the directional alignment between embeddings, effectively capturing semantic similarities irrespective of their magnitudes.
The optimization process utilizes the InfoNCE loss function [57], defined as:
where is the temperature parameter that scales the cosine similarities to control the distribution’s concentration. This approach ensures that the model effectively learns to distinguish relevant from irrelevant information across different modalities, thereby enhancing its performance in multimodal retrieval tasks.
Hard Negatives
The quality and diversity of negative samples are essential for improving contrastive learning [53]. Inspired by ANCE [69], we employ a two-stage training strategy: (1) Initial Training: We first train the model using randomly selected negative candidates, resulting in Model . (2) Hard Negative Mining and Continue Training: Using , we retrieve the top candidates for each query and select non-relevant candidates from them as hard negatives. We then use these hard negatives to further train , refining it into the final model. This ensures that the model can learn from both easily distinguishable and more challenging examples, thereby enhancing performance.
Training Data Composition
A critical factor in multimodal representation learning is the composition of training data. Although previous studies like [22] have demonstrated that MLLMs can develop multimodal representation capabilities after being fine-tuned on single-modal data, the effect of data diversity on model performance remains unclear. Therefore, we compare the performance of models trained with different data combinations across various retrieval scenarios within our classification principle. Specifically, we used four types of training data: single-modal (including TT and II), cross-modal (including TVD and TI), fused-modal training data (including ITIT), and a mixed dataset combining the first three types. These different training data types result in a total of six models.
For single-modal data, we utilized the TT dataset from MSMARCO [47] and the II dataset from ImageNet [10], treating images within the same category as positive matches and those from different categories as negatives. For cross-modal data, we employed TI pairs from the LAION [54] dataset and TVD pairs from the Docmatix [27] dataset. For fused-modal data, we use the EVQA [46] dataset (ITIT). For each subcategory, we randomly sampled 100,000 training instances to train the models independently. For the mixed dataset, we uniformly sampled 20,000 instances from each of the five datasets to train the final model, ensuring fair and reliable comparative experimental results. The performance of these six models on the UMRB-Partial test dataset is presented in Figure 3.
The results indicate that: (1) Models trained on single data types excel in corresponding retrieval tasks. For instance, models trained on TT data performed best in text retrieval tasks.333Detail results are shown in the Appendix Table 8. (2) A balanced mix of different data types enhanced performance across various settings. This suggests that increasing the diversity of training modalities effectively improves the model’s overall retrieval capabilities.
The above analysis highlights the importance of adequately representing each data type in training datasets to develop models that meet the requirements of universal multi-modal retrieval. During data collection, we observed that single-modal and cross-modal data are abundant, with over ten million training instances available. In contrast, fused-modal data remains limited. Common fused-modal training datasets such as EVQA[46], INFOSEEK[6], and CIRR [38] collectively contain fewer than one million instances. Additionally, these existing fused-modal datasets cover only a limited range of domains. Thus, efficiently supplementing high-quality fused-modal training data is essential. To address this challenge, we propose leveraging the generative capabilities of LLMs and MLLMs to synthesize additional training data.
4.2 Fused-Modal Data Synthesis
To efficiently synthesize high-quality data while minimizing manual intervention, we adopt a strategy similar to Doc2Query [15]. However, our approach differs in that we aim to generate fuse-modal candidate-to-query relevance data instead of single-modality, text-based relevance pairs. This requires obtaining high-quality candidates that include both image and text content. We primarily extracted such data from Wikipedia paragraphs444github.com/google-research-datasets/wit/blob/main/wikiweb2m.md. Additionally, to enhance the domain diversity of the candidate data, we employed a domain classification model555hf.co/facebook/bart-large-mnlifacebook to perform fine-grained classification of Wikipedia data into categories such as animals and plants. We then uniformly sampled from these categories and retained data with classification confidence scores above 0.5. Ultimately, we obtained 313,284 candidate entries, each containing both text and image content.
Based on the prepared data, the overall synthesis pipeline (Figure 4) could be divided into the following steps:
Doc2Query Generation: The passage content from each candidate is input into an LLM666In the entire pipeline, we utilize Qwen2.5-72B-Instruct as our LLM. using a prompt to generate a natural query. To ensure the quality of the generated queries, we built a vector index of all passage contents using a text vector retrieval model777hf.co/Alibaba-NLP/gte-Qwen2-1.5B-instruct. Each generated query is then used to retrieve the corresponding passage from this collection. If the passage associated with the query is not within the top 20 retrieved items, the query is considered low quality due to low relevance and is discarded. In this step, we discarded 1.2% of the total generated queries. This process allows us to construct TIT training data.
Entity Extraction and Query Rewrite: We aim for the synthesized queries to include both texts and images (i.e., ITIT type). To achieve this, we leverage entity extraction followed by image retrieval for the extracted entities and caption generation to supplement the image data on the query side. Specifically, for each generated query from the first step, we prompt the LLM to extract entities from it with the text passage as reference, and then rewrite the original query into . For example, the query “Where is Iris pseudacorus native?” is transformed by the model to the rewritten query ”Where is the native habitat of this plant?” with the entity ”Iris pseudacorus” extracted. We then seek images that match this entity and combine them with the rewritten query to form the final fuse-modal query.
Image Retrieval and Generation: We explore two methods for obtaining images. The first method uses the Google Image Search API888https://serpapi.com/google-images-api to retrieve images matching the entity terms, retaining the top five results. The second method involves generating images using a text-to-image model999 https://hf.co/black-forest-labs/FLUX.1-dev. Specifically, we first use the LLM to generate a caption suitable for image generation based on the entity and the passage of the generated query, then input this caption into the text-to-image generation model to create the corresponding image. This approach allows us to quickly and efficiently obtain high-quality, diverse images. The synthesized results can also be assembled into IT IT retrieval type data.
Data Filtering: To ensure the quality of the synthesized data, we perform filtering [9] on the final dataset. We observe that images generated by the FLUX model have consistent quality, whereas images retrieved via the Google Image Search API often include noisy data. Therefore, for images obtained through the Google Image Search API, we use the CLIP model101010https://hf.co/openai/clip-vit-large-patch14 to assess image-caption relevance. Images with a relevance score below were filtered out.
Through the synthesis pipeline, we produce 1,135,000 high-quality fuse-modal training data entries (including TIT and ITIT types). After filtering, we retain 1,102,000 entries, resulting in a data loss rate of . The entire process consumed 600 A100 GPU hours. Detailed descriptions of all prompts used in the data synthesis pipeline and examples of the synthesized data are provided in the Appendix §10.
| UMRB | Size | Single-Modal | Cross-Modal | Fused-Modal | Avg. | ||||||
| Task (#Datasets) | TT (16) | II (1) | TI (4) | TVD (10) | IT (4) | TIT (2) | ITT (5) | ITI (2) | ITIT (3) | (47) | |
| VISTA [74] | 0.2B | 55.15 | 31.98 | 32.88 | 10.12 | 31.23 | 45.81 | 53.32 | 8.97 | 26.26 | 37.32 |
| CLIP-SF [66] | 0.4B | 39.75 | 31.42 | 59.05 | 24.09 | 62.95 | 66.41 | 53.32 | 34.90 | 55.65 | 43.66 |
| One-Peace [64] | 4B | 43.54 | 31.27 | 61.38 | 42.9 | 65.59 | 42.72 | 28.29 | 6.73 | 23.41 | 42.01 |
| DSE [41] | 4.2B | 48.94 | 27.92 | 40.75 | 78.21 | 52.54 | 49.62 | 35.44 | 8.36 | 40.18 | 50.04 |
| E5-V [22] | 8.4B | 52.41 | 27.36 | 46.56 | 41.22 | 47.95 | 54.13 | 32.90 | 23.17 | 7.23 | 42.52 |
| GME-Qwen2VL-2B | 2.2B | 55.93 | 29.86 | 57.36 | 87.84 | 61.93 | 76.47 | 64.58 | 37.02 | 66.47 | 64.45 |
| GME-Qwen2VL-7B | 8.2B | 58.19 | 31.89 | 61.35 | 89.92 | 65.83 | 80.94 | 66.18 | 42.56 | 73.62 | 67.44 |
5 Experiments
5.1 Settings
Training Data
Building on the findings from §4.1, we train our model using a diverse dataset of 8 million instances spanning various retrieval modalities. For single-modal retrieval tasks, we utilize datasets including MSMARCO [47], NQ [26], HotpotQA [70], TriviaQA [24], SQuAD [52], FEVER [56], and AllNLI for SimCSE [14], selecting a total of 1 million entries. From ImageNet [10], we extract 1 million image-to-image training instances, designating images within the same class as positive samples and others as negative samples. For cross-modal retrieval tasks, we incorporate 2 million entries from the LAION [54], MSCOCO [32], and Docmatix [27] datasets. Additionally, for fused-modal retrieval tasks, we include a total of 2 million instances: million synthesized by us, and the remaining from the M-BEIR [66] training data.
Training Configuration
We use Qwen2-VL [65] model series as the backbone for our MLLM, conducting training on models with both 2 billion (2B) and 7 billion (7B) parameters. Our training utilizes Low-Rank Adaptation (LoRA) [19] with a rank of 8, a learning rate of 1e-4, and a temperature setting of 0.03. To manage the varying number of visual tokens required by Qwen2-VL for different image resolutions and maintain training efficiency, we limit the maximum number of visual tokens per image to 1,024.
For data with images, we set the maximum text length to 1,800 tokens, using a batch size of 128 for the 2B model and 32 for the 7B model. For text-only data, the maximum length was set to 512 tokens, with batch size of 512 for the 2B model and 128 for the 7B model. Each training sample included 8 negative examples. To conserve GPU memory, we employ gradient checkpointing [5] and train the model using bfloat16 precision. All training was conducted on eight NVIDIA A100 GPUs, each with 80GB of memory.
Baselines
We compare our method against four types of retrieval systems: (1) Previous representative UMR models, for example, VISTA [74] for text encoder based, and E5-V [22] for MLLM based; (2) Powerful multimodal representation (embedding) models, i.e. One-Peace [64], which supports modalities beyond text and image and hence could also be tested on our UMRB; (3) Recent visual document retrieval models, namely DSE [41]; and (4) the classic cross-modal pipeline, CLIP score-fusion, denoted as CLIP-SF, which provides top-tier cross-modal performance. We exclude comparisons with state-of-the-art text retrieval models as VISTA demonstrates comparable performance levels.
5.2 Main Results
Table 3 presents the evaluation results of the baseline systems alongside our proposed GME. Scores are averaged across each sub-task and categorized by retrieval modality type: single-modal, cross-modal, and fused-modal. Additionally, the overall micro-average score on the UMRB is in the last column. First, focusing on the average scores, our smaller model, i.e. GME-Qwen2-VL-2B, already outperforms the previous state-of-the-art UMR model (VISTA [74]). The larger model, i.e. GME-Qwen2-VL-7B, further enhances this performance, demonstrating the effectiveness of our approach in handling UMR tasks.
Second, our models outperform smaller methods such as VISTA (million-level parameters) and One-Peace (4B parameters). The larger MLLM baseline, E5-V [22] (8B parameters), performs well in text-dominated tasks (e.g., TT) but falls short in other areas. This indicates that training with multimodal data is crucial for achieving superior performance in UMR tasks. Our training data provides a stronger foundation for future advancements.
Next, the cross-modal pipeline CLIP-SF outperforms UMR models like VISTA, E5-V, and One-Peace. For VISTA and E5-V, the performance gap is likely due to limitations in their text-modality bounds: VISTA is constrained by the text embedding space of its fixed backbone, and E5-V is limited by text-only training. One-Peace’s modality alignment-centered modeling may not be optimized for fused-modal content. In contrast, our models are specifically designed to handle fused-modal data, resulting in significantly better performance compared to the baselines. Although our training data includes several previously constructed fused-modal datasets, the contribution of our generated fused-modal training data will be discussed in §5.3.
Finally, we compare with the recent visual document retrieval model DSE [41], specialized for the TVD task within the Cross-Modal group, which has approximately 4B parameters. Our models are competitive with or exceed the performance of this task-specific baseline, demonstrating the feasibility and promise of integrating visual document retrieval into a unified retriever framework.
5.3 Analyses
Are the Produced Embeddings Modality Universal?
Given our the impressive performance of our model, we assess the quality of its embeddings. Specifically, we investigate whether the embeddings are modality-universal meaning that embeddings representing the same semantic content across different modalities are closely clustered in the embedding space, or if they remain in separate sub-spaces tailored for each modality-specific task. To probe this question, we sample 1000 instances from the EVQA dataset and visualize their embeddings of different modalities by t-SNE, as shown in Figure 5. We also highlight two semantic close groups with yellow and pink labels, respectively. We can observe that the embeddings from CLIP are distinctly separated by modality, whereas the embeddings from our model are intermingled and organized semantically. Meanwhile, the points from the same semantic group are closely clustered. This demonstrates that our model effectively generates modality-universal representations, enhancing its applicability across various UMR tasks.
| Setting | Single | Cross | Fused | Average |
| w/ EVQA | 45.13 | 60.21 | 49.32 | 51.55 |
| w/ Gen | 46.27 | 61.19 | 51.46 | 52.97 |
| w/ Gen | 47.08 | 61.35 | 52.01 | 53.48 |
Ablation Study on Synthetic Fused-Modal Data
We propose an efficient data synthesis pipeline (§4.2) and generate large-scale fused-modal pairs to support model training. After witnessing the state-of-the-art performance of our model, it is natural to question the contribution of this synthetic data to the overall performance. To this end, we conduct an ablation study using three parallel training datasets, each comprising 100,000 pairs: original EVQA data, synthetic data with Google-retrieved images (GenGoogle), and synthetic data with FLUX-generated images (GenFlux). We train three models with identical parameters on these datasets and evaluate their performance on UMRB-Partial, with results shown in Table 4. Both synthetic datasets outperform the original EVQA data, indicating the high quality of our synthesized data. Although Google-retrieved images achieved marginally better performance than FLUX-generated images, the difference is minor and acceptable given the potential limitations of the Google Search API for rapid, large-scale dataset generation.
Training Scaling Law
Our approach is primarily data-centric, constructing a diverse training dataset of approximately million samples across various UMR settings (§5.1). Training on such a large-scale dataset demands significant computational resources and time. Therefore, we explored the training scaling law by examining how model performance evolves with increasing training steps. Due to the time-consuming nature of evaluating certain retrieval tasks, we assessed performance on our UMRB-Partial dataset for faster evaluation. Figure 6 illustrates the performance progression of our 2B and 7B models on UMRB-Partial during training. Both models exhibit linear performance improvements as training continues, suggesting that extended training could yield further benefits. However, due to time constraints, we halted current training. Future work will investigate longer training periods to enhance model performance further.
Ablation Study on Modeling
We conduct an ablation study to investigate the effectiveness of different design choices of GME. We consider the following three aspects: (1) Fine-tuning strategy. Our final models are trained by LoRA with rank 8. We compare with other rank values and full fine-tuning. The results in the first group of Table 5 show that LoRA with rank 8 yields the best performance. (2) Training data organization. We compare models trained without hard negative mining. The second group of Table 5 demonstrates that the removal of hard negatives led to performance declines, indicating that it is essential for effective retrieval model training. (3) Retrieval instructions. We compare models trained without retrieval instructions. The third group shows that retrieval instructions are crucial for better UMR. (4) Modeling techniques. Our final models are in the casual attention mode and use the EOS token state as the embedding, hence we compare the performance of the model trained with mean pooling and the bi-directional attention mechanism. The last group of Table 5 shows that these alternative settings negatively impact performance.
| Setting | Single | Cross | Fused | Average |
| Fine-tuning strategy | ||||
| LoRA r=8 | 48.09 | 78.39 | 51.88 | 59.45 |
| LoRA r=16 | 47.86 | 78.63 | 51.42 | 59.30 |
| LoRA r=32 | 47.85 | 78.55 | 50.48 | 58.96 |
| LoRA r=64 | 47.65 | 78.61 | 51.09 | 59.11 |
| Full training | 43.16 | 75.79 | 49.28 | 56.07 |
| Training data organization | ||||
| w/o hard-negative | 47.55 | 78.01 | 50.95 | 58.83 |
| Retrieval Setting | ||||
| w/o Instruction | 46.82 | 78.10 | 49.09 | 58.00 |
| Model Design | ||||
| w/ mean pooling | 47.86 | 77.95 | 51.33 | 59.04 |
| w/ bi-attention | 46.55 | 76.78 | 49.54 | 57.62 |
6 Conclusion
In this work, we target the universal multimodal retrieval (UMR) problem. We begin by systematically categorizing current UMR tasks, proposing a comprehensive classification framework. Based on this, we explore ways to further improve MLLM-based UMR models, suggesting the GME model. The GME models are trained using contrastive learning loss on a diverse set of multimodal data settings, while also extending support for visual retrieval. Additionally, to overcome limitations in existing UMR evaluation benchmarks, we compiled a new comprehensive benchmark (i.e., UMRB) by integrating multiple data sources. This benchmark effectively balances existing UMR tasks with the increasingly important text and visual document retrieval tasks, enabling a more thorough assessment of UMR model performance. We evaluate existing UMR models and our proposed GME model on UMRB, finding that our model achieves state-of-the-art performance. We also conducted various analyses to validate the effectiveness of our methods and enhance our understanding of them. Our benchmark, models, and other materials are open-source at https://hf.co/Alibaba-NLP/gme-Qwen2-VL-7B-Instruct.
Acknowledgments
This work receives partial support from the Natural Science Foundation of China (under Grant 624B2048) and Research Grant Council of Hong Kong (PolyU/15209724).
Appendix
7 UMRB Details
| Name | Type | Categ. | Eval | Candidates | Eval Query | Eval Candidate | In partial |
| Samples | Nums | avg. chars | avg. chars | ||||
| ArguAna | Single-Modal | TT | 10,080 | 1,406 | 192.98 | 166.80 | True |
| Climate-FEVER | Single-Modal | TT | 1,535 | 5,416,593 | 20.13 | 84.76 | False |
| CQADupStack | Single-Modal | TT | 13,145 | 457,199 | 8.59 | 129.09 | False |
| DBPedia | Single-Modal | TT | 400 | 4,635,922 | 5.39 | 49.68 | False |
| FEVER | Single-Modal | TT | 6,666 | 5,416,568 | 8.13 | 84.76 | False |
| FiQA2018 | Single-Modal | TT | 648 | 57,638 | 10.77 | 132.32 | False |
| HotpotQA | Single-Modal | TT | 7,405 | 5,233,329 | 17.61 | 46.30 | False |
| MSMARCO | Single-Modal | TT | 6,980 | 8,841,823 | 5.96 | 55.98 | False |
| NFCorpus | Single-Modal | TT | 323 | 3,633 | 3.30 | 232.26 | True |
| NQ | Single-Modal | TT | 3,452 | 2,681,468 | 9.16 | 78.88 | False |
| Quora | Single-Modal | TT | 10,000 | 522,931 | 9.53 | 11.44 | True |
| SCIDOCS | Single-Modal | TT | 1,000 | 25,657 | 9.38 | 176.19 | True |
| SciFact | Single-Modal | TT | 300 | 5,183 | 12.37 | 213.63 | False |
| Touche2020 | Single-Modal | TT | 49 | 382,545 | 6.55 | 292.37 | False |
| TRECCOVID | Single-Modal | TT | 50 | 171,332 | 10.60 | 160.77 | True |
| WebQA | Single-Modal | TT | 2,455 | 544,457 | 18.58 | 37.67 | False |
| Nights | Single-Modal | II | 2,120 | 40,038 | - | - | True |
| VisualNews | Cross-Modal | TI | 19,995 | 542,246 | 18.78 | - | False |
| Fashion200k | Cross-Modal | TI | 1,719 | 201,824 | 4.89 | - | False |
| MSCOCO | Cross-Modal | TI | 24,809 | 5,000 | 10.43 | - | True |
| Flickr30k | Cross-Modal | TI | 5,000 | 1,000 | 12.33 | - | True |
| TAT-DQA | Cross-Modal | TVD | 1,646 | 277 | 12.44 | - | False |
| ArxivQA | Cross-Modal | TVD | 500 | 500 | 17.12 | - | False |
| DocVQA | Cross-Modal | TVD | 451 | 500 | 8.23 | - | True |
| InfoVQA | Cross-Modal | TVD | 494 | 500 | 11.29 | - | False |
| Shift Project | Cross-Modal | TVD | 100 | 1,000 | 16.01 | - | True |
| Artificial Intelligence | Cross-Modal | TVD | 100 | 968 | 12.3 | - | False |
| Government Reports | Cross-Modal | TVD | 100 | 972 | 12.62 | - | False |
| Healthcare Industry | Cross-Modal | TVD | 100 | 965 | 12.56 | - | False |
| Energy | Cross-Modal | TVD | 100 | 977 | 13.49 | - | False |
| TabFQuad | Cross-Modal | TVD | 280 | 70 | 16.49 | - | False |
| VisualNews | Cross-Modal | IT | 20,000 | 537,568 | - | 18.53 | False |
| Fashion200k | Cross-Modal | IT | 4,889 | 61,707 | - | 4.95 | False |
| MSCOCO | Cross-Modal | IT | 5,000 | 24,809 | - | 10.43 | True |
| Flickr30k | Cross-Modal | IT | 1,000 | 5,000 | - | 12.33 | True |
| WebQA | Fused-Modal | TIT | 2,511 | 403,196 | 16.43 | 12.83 | False |
| EDIS | Fused-Modal | TIT | 3,241 | 1,047,067 | 20.07 | 15.53 | False |
| OVEN | Fused-Modal | ITT | 50,004 | 676,667 | 6.52 | 82.13 | False |
| INFOSEEK | Fused-Modal | ITT | 11,323 | 611,651 | 8.76 | 91.49 | False |
| ReMuQ | Fused-Modal | ITT | 3,609 | 138,794 | 13.82 | 34.26 | True |
| OKVQA | Fused-Modal | ITT | 5,046 | 114,516 | 8.09 | 102.55 | True |
| LLaVA | Fused-Modal | ITT | 5,120 | 5,994 | 10.70 | 90.65 | True |
| FashionIQ | Fused-Modal | ITI | 6,003 | 74,381 | 11.70 | - | True |
| CIRR | Fused-Modal | ITI | 4,170 | 21,551 | 11.01 | - | True |
| OVEN | Fused-Modal | ITIT | 14,741 | 335,135 | 5.91 | 94.76 | True |
| EVQA | Fused-Modal | ITIT | 3,743 | 68,313 | 9.38 | 211.12 | False |
| INFOSEEK | Fused-Modal | ITIT | 17,593 | 481,782 | 7.94 | 96.00 | False |
Table 6 summarizes all UMRB tasks along with their statistics. Table 14 provides examples of different task types. Below is a brief description of each dataset included in the UMRB.
7.1 Single-Modal Tasks
WebQA [4]
This dataset is derived from Wikipedia. In the TT setup, both the query and candidate are text. The objective is to find a Wikipedia paragraph that answers the question. We have used 2,455 samples as the test set.
Nights [13]
This dataset contains human judgments on the similarity of various image pairs, where both the query and candidate are images. The task is to identify an image that resembles the provided query image. We included 2,120 samples in our UMRB.
ArguAna, ClimateFEVER, CQADupstack, DBPedia, FEVER, FiQA2018, HotpotQA, MSMARCO, NFCorpus, NQ, Quora, SCIDOCS, SciFact, Touche2020 and TRECCOVID
For these datasets, we use the processed versions from BEIR [55].
7.2 Cross-Modal Tasks
VisualNews [34]
This dataset focuses on the news domain and consists of pairs of news headlines and associated images. In UMRB, this dataset can be transformed into two tasks: retrieving the corresponding image based on the news headline (TI) and retrieving the corresponding news headline based on the image (IT). We utilized 19,995 and 20,000 samples to construct the test set.
Fashion200k [16]
This dataset includes pairs of images and product descriptions. In total, we have 1,719 instances for the task TI and 4,889 instances for the task IT for evaluation.
MSCOCO [32]
This dataset is a well-known image caption dataset. Similar to VisualNews, it is converted into two tasks: “IT”, which retrieves the caption given an image and “TI”, which retrieves the image given a caption.
Flickr30k[50]
This dataset consists of images paired with detailed textual descriptions. We have a total of 1,000 instances for the IT task and 5,000 instances for the TI task available for evaluation.
TAT-DQA, ArxivQA, DocVQA, InfoVQA, Shift Project, Artificial Intelligence, Government Reports, Healthcare Industry, Energy, TabFQuad
These datasets constitute the retrieval task of TVD. Their queries are standard questions, and the candidates are document screenshots. For these datasets, we used the processed versions from ViDoRe [12].
7.3 Fused-Modal Tasks
WebQA [4]
Similar to WebQA in the Single-Modal setting, this dataset is also derived from Wikipedia, but in the TIT setup, the candidates consist of images and text. The task is to find a Wikipedia paragraph with accompanying text and images to answer a specific question. There are 2,511 samples in the evaluation set.
EDIS [37]
This dataset involves the cross-modal image search within the news domain. The queries are texts containing entities and events, with candidates consisting of news images and their accompanying headlines. The task requires the model to comprehend both entities and events from the text queries and retrieve the corresponding image and headline.
OVEN [20]
The dataset is sourced from Wikipedia, where a query consists of an image and a question related to the image. The candidates are the Wikipedia title along with the first 100 tokens of its summary. If the associated Wikipedia content includes images, it constitutes an ITIT task; otherwise, it forms an ITT task. In the evaluation, we have 14,741 samples for the ITIT task and 50,004 samples for the ITT task.
INFOSEEK [6]
This dataset is similar to OVEN, with queries consisting of images alongside text questions. The candidates are Wikipedia snippets of 100 tokens containing the exact answers. This dataset also encompasses two tasks: for the ITIT and ITT tasks, we used 17,593 and 11,323 samples, respectively.
ReMuQ [40]
The dataset is augmented from the WebQA questions by adding images to create new multimodal queries along with a large text corpus. For evaluation, we used 3,609 instances from this dataset.
OKVQA [43]
This dataset includes visual questions that require external knowledge to answer. It is structured as an ITT retrieval task, where queries consist of visual questions containing images and text, with candidates being external knowledge sources that can assist in answering the questions.
LLaVA [33]
This dataset contains high-quality conversations about an image generated by GPT-3.5, involving exchanges between a human and an AI assistant. The queries comprise questions and instructions sent by humans to the AI assistant, which include both images and text, while the candidates are the AI assistant’s replies. We utilized 5,120 samples from this dataset in the UMRB evaluation.
FashionIQ [67]
This dataset features images of fashion products along with crowd-sourced descriptions that highlight the differences between these products. Each query consists of an image and a modification sentence that describes changes to the given image, with the retrieval target being the specified image. In the UMRB evaluation, we used 6,003 samples from this dataset.
CIRR [38]
Similar to FashionIQ, CIRR can also be used for composed image retrieval. It involves pairs of real-life reference and target images in each test case, along with a modification sentence detailing the differences between the two images. For the UMRB evaluation, we utilized 4,170 samples from this dataset.
EVQA [46]
This dataset is akin to INFOSEEK, with the key distinction being that the retrieval target of EVQA is a complete Wikipedia paragraph with a maximum length of several thousand tokens. We used 3,743 samples for evaluation, eliminating multi-hop issues present in the original test set. We selected Wikipedia paragraphs from the original dataset as candidates and supplemented them with images. Images native to each paragraph were included when available; otherwise, the first image from the article was utilized due to its typically representative nature.
7.4 UMRB-Partial
The full UMRB dataset consists of 47 subtasks, approximately 200,000 evaluation instances, and 40 million candidates, resulting in a significant overhead when testing the model. During our experiments with the GME-7B model, a full evaluation required approximately 400 A100*80G GPU hours. To facilitate development and verification, we created a smaller benchmark by condensing the complete UMRB, which we refer to as UMRB-Partial. Column 8 of Table 6 indicates whether a dataset is included in UMRB-Partial. Testing the GME-7B model on UMRB-Partial reduced the evaluation time from 400 A100*80G GPU hours to 80 A100*80G GPU hours.
8 Results Details
In this section, we present the detailed scores achieved by our GME and the baseline models on various tasks. Additionally, we provide results from other benchmarks, including BEIR, M-BEIR, and ViDoRe.
| Type | Task | Dataset | VISTA | CLIP-SF | One-Peace | DSE | E5-V | GME-2B | GME-7B |
| Single- Modal | TT | ArguAna | 63.61 | 52.45 | 32.93 | 53.46 | 54.28 | 63.18 | 72.11 |
| Climate-FEVER | 31.17 | 20.00 | 20.27 | 19.79 | 21.64 | 41.08 | 48.36 | ||
| CQADupStack | 42.35 | 30.61 | 41.32 | 36.51 | 41.69 | 39.06 | 42.16 | ||
| DBPedia | 40.77 | 26.37 | 32.43 | 40.75 | 38.78 | 41.00 | 46.30 | ||
| FEVER | 86.29 | 50.58 | 51.91 | 80.12 | 78.99 | 92.06 | 93.81 | ||
| FiQA2018 | 40.65 | 22.14 | 36.79 | 36.2 | 45.41 | 43.8 | 63.23 | ||
| HotpotQA | 72.6 | 41.33 | 46.51 | 70.79 | 60.88 | 65.3 | 68.18 | ||
| MSMARCO | 41.35 | 22.15 | 36.55 | 37.73 | 41.23 | 40.61 | 42.93 | ||
| NFCorpus | 37.39 | 27.05 | 31.6 | 32.82 | 36.97 | 38.84 | 36.95 | ||
| NQ | 54.15 | 25.45 | 42.87 | 52.97 | 51.58 | 54.52 | 56.08 | ||
| Quora | 88.90 | 81.63 | 87.46 | 85.84 | 87.6 | 88.12 | 89.67 | ||
| SCIDOCS | 21.73 | 14.75 | 21.64 | 15.66 | 22.36 | 22.94 | 26.35 | ||
| SciFact | 74.04 | 55.98 | 64.51 | 68.97 | 72.75 | 74.19 | 82.43 | ||
| Touche2020 | 25.7 | 17.47 | 16.90 | 14.50 | 21.61 | 26.57 | 22.55 | ||
| TRECCOVID | 77.90 | 63.61 | 69.28 | 52.98 | 72.85 | 71.73 | 77.49 | ||
| WebQA | 83.80 | 84.44 | 63.67 | 83.95 | 89.94 | 94.34 | 94.34 | ||
| II | Nights | 24.43 | 31.42 | 31.27 | 27.36 | 27.92 | 30.61 | 30.57 | |
| Cross- Modal | TI | VisualNews | 5.77 | 42.80 | 48.95 | 14.12 | 29.46 | 39.20 | 46.27 |
| Fashion200k | 3.08 | 18.38 | 32.34 | 3.08 | 3.78 | 23.50 | 27.64 | ||
| MSCOCO | 47.97 | 80.75 | 71.45 | 74.62 | 52.38 | 76.22 | 79.77 | ||
| Flickr30k | 74.68 | 94.28 | 92.78 | 94.42 | 77.38 | 94.5 | 97.38 | ||
| TVD | TAT-DQA | 2.05 | 5.49 | 14.44 | 49.01 | 9.08 | 57.88 | 64.06 | |
| ArxivQA | 10.30 | 24.10 | 43.94 | 78.17 | 41.16 | 81.41 | 82.55 | ||
| DocVQA | 8.01 | 11.80 | 23.48 | 45.83 | 24.37 | 46.86 | 49.34 | ||
| InfoVQA | 30.02 | 48.78 | 59.97 | 82.06 | 49.5 | 84.97 | 88.79 | ||
| Shift Project | 3.26 | 6.06 | 17.02 | 69.84 | 13.16 | 77.94 | 83.5 | ||
| Artificial Intelligence | 7.34 | 28.64 | 45.41 | 96.88 | 46.18 | 95.75 | 98.02 | ||
| Government Reports | 6.90 | 34.67 | 55.98 | 92.04 | 53.05 | 92.05 | 94.05 | ||
| Healthcare Industry | 9.39 | 32.64 | 59.55 | 96.35 | 59.61 | 96.08 | 97.29 | ||
| Energy | 11.05 | 27.19 | 53.21 | 92.62 | 56.77 | 89.17 | 93.09 | ||
| TabFQuad | 13.08 | 21.53 | 57.05 | 79.29 | 58.22 | 91.79 | 94.92 | ||
| IT | VisualNews | 2.79 | 42.67 | 47.27 | 8.74 | 29.54 | 38.21 | 47.16 | |
| Fashion200k | 4.72 | 18.10 | 30.89 | 3.91 | 4.62 | 26.61 | 31.05 | ||
| MSCOCO | 48.92 | 91.94 | 85.6 | 82.06 | 86.4 | 85.18 | 85.92 | ||
| Flickr30k | 68.50 | 99.11 | 98.60 | 97.11 | 89.62 | 99.00 | 98.9 | ||
| Fused- Modal | TIT | WebQA | 54.84 | 78.42 | 32.42 | 66.99 | 49.62 | 82.24 | 84.11 |
| EDIS | 36.78 | 54.09 | 53.01 | 41.26 | 49.62 | 68.10 | 77.40 | ||
| ITT | OVEN | 22.32 | 45.98 | 23.69 | 0.38 | 14.4 | 59.67 | 64.13 | |
| INFOSEEK | 18.53 | 27.58 | 20.05 | 3.06 | 12.69 | 39.22 | 34.67 | ||
| ReMuQ | 76.20 | 83.71 | 26.41 | 94.60 | 52.15 | 96.73 | 95.48 | ||
| OKVQA | 17.14 | 17.44 | 9.67 | 13.28 | 16.71 | 30.08 | 32.61 | ||
| LLaVA | 72.81 | 91.91 | 51.64 | 53.18 | 77.48 | 98.93 | 98.18 | ||
| ITI | FashionIQ | 3.28 | 24.54 | 2.93 | 9.81 | 3.73 | 26.34 | 29.89 | |
| CIRR | 14.65 | 45.25 | 10.53 | 36.52 | 13.19 | 47.70 | 51.79 | ||
| ITIT | OVEN | 27.77 | 68.83 | 30.56 | 0.39 | 54.46 | 78.96 | 83.05 | |
| EVQA | 28.75 | 40.08 | 16.64 | 15.34 | 26.39 | 77.32 | 79.88 | ||
| INFOSEEK | 22.27 | 49.05 | 23.32 | 5.96 | 39.69 | 41.14 | 31.58 | ||
| Avg. | 37.32 | 43.66 | 42.01 | 50.04 | 42.52 | 63.42 | 65.87 |
8.1 Detailed Results on UMRB
Table 7 presents the detailed evaluation results of the baseline systems alongside our GME on UMRB tasks. First, focusing on the average scores, our smaller model, i.e. GME-Qwen2-VL-2B, already outperforms the previous state-of-the-art UMR model (VISTA). The larger model, i.e. GME-Qwen2-VL-7B, further enhances this performance. In addition, focusing on specific scores on different datasets, our GME achieves state-of-the-art performance on each dataset except the Nights dataset. VISTA and CLIP-SF scored highly on the Nights dataset, likely due to their use of independent image and text encoders for cross-modal retrieval. In the II task, these models relied solely on the image encoder for encoding without cross-modal alignment, which may explain their superior performance on the Nights dataset.
| Type | Task | Dataset | TT | II | TVD | TI | ITIT | Mix |
| Single- Modal | TT | Arguan | 56.25 | 43.51 | 56.73 | 33.53 | 53.22 | 56.22 |
| NFCorpus | 35.23 | 28.89 | 33.23 | 33.18 | 30.48 | 35.76 | ||
| Quora | 87.82 | 74.37 | 86.32 | 86.43 | 85.2 | 87.4 | ||
| SCIDOCS | 19.07 | 11.82 | 17.51 | 17.2 | 16.93 | 19.88 | ||
| TRECCOVID | 75.57 | 47.89 | 50.89 | 72.37 | 58.92 | 76.38 | ||
| II | Nights | 27.97 | 28.11 | 24.9 | 28.53 | 26.04 | 30.85 | |
| Cross- Modal | TI | MSCOCO | 59.7 | 59.41 | 63.67 | 76.91 | 44.97 | 75.3 |
| Flickr30k | 83.92 | 65.52 | 87.32 | 93.18 | 74.52 | 93.06 | ||
| TVD | DocVQA | 35.8 | 24.24 | 48.38 | 40.58 | 28.05 | 45.62 | |
| Shift Project | 57.86 | 45.47 | 77.08 | 50.36 | 53.12 | 74.84 | ||
| IT | MSCOCO | 74.72 | 63.82 | 80.46 | 84.64 | 70.48 | 84.24 | |
| Flickr30k | 94.1 | 82.5 | 96.3 | 97.2 | 90.1 | 97.5 | ||
| Fused- Modal | ITT | LLaVA | 92.75 | 89.05 | 86.02 | 89.24 | 88.73 | 95.02 |
| ReMuQ | 89.61 | 85.47 | 76.45 | 85.12 | 86.73 | 89.75 | ||
| OKVQA | 24.55 | 16.6 | 15.78 | 16.92 | 18.57 | 20.23 | ||
| ITI | FashionIQ | 5.53 | 4.2 | 5.43 | 8.86 | 11.08 | 11.89 | |
| CIRR | 17.24 | 15.04 | 15.42 | 17.5 | 25.71 | 29.86 | ||
| ITIT | OVEN | 59.81 | 38.42 | 57.31 | 56.69 | 65.08 | 63.04 | |
| Avg. | 55.42 | 45.80 | 54.50 | 54.91 | 51.55 | 60.38 |
8.2 Detailed Results on UMRB-Partial
Figure 3 of main paper illustrates our exploration of the training data, as discussed in Section 4.2, with specific results presented in Table 8. This table details the scores of our models trained on six data types: TT, II, TVD, TI, ITIT, and Mix across various tasks. We find that the model trained on mixed data performs the best.
| BEIR | Avg. |
Fever | FiQA |
NQ | Quora |
| Text Embedder | ||||||||||||||||||||||||||||||||||||||||||
| gte-Qwen2-7B-instruct | 60.25 | 64.27 | 45.88 | 46.43 | 52.42 | 95.11 | 62.03 | 73.08 | 45.98 | 40.6 | 67 | 90.09 | 28.91 | 79.06 | 30.57 | 82.26 | ||||||||||||||||||||||||||
| NV-Embed-v1 | 59.36 | 68.2 | 34.72 | 50.51 | 48.29 | 87.77 | 63.1 | 79.92 | 46.49 | 38.04 | 71.22 | 89.21 | 20.19 | 78.43 | 28.38 | 85.88 | ||||||||||||||||||||||||||
| gte-Qwen2-1.5B-instruct | 58.29 | 69.72 | 42.91 | 44.76 | 48.69 | 91.57 | 54.7 | 68.95 | 43.36 | 39.34 | 64 | 89.64 | 24.98 | 78.44 | 27.89 | 85.38 | ||||||||||||||||||||||||||
| voyage-large-2-instruct | 58.28 | 64.06 | 32.65 | 46.6 | 46.03 | 91.47 | 59.76 | 70.86 | 40.6 | 40.32 | 65.92 | 87.4 | 24.32 | 79.99 | 39.16 | 85.07 | ||||||||||||||||||||||||||
| neural-embedding-v1 | 58.12 | 67.21 | 32.3 | 49.11 | 48.05 | 89.46 | 58.94 | 78.87 | 42 | 42.6 | 68.36 | 89.02 | 27.69 | 78.82 | 24.06 | 75.33 | ||||||||||||||||||||||||||
| GritLM-7B | 57.41 | 63.24 | 30.91 | 49.42 | 46.6 | 82.74 | 59.95 | 79.4 | 41.96 | 40.89 | 70.3 | 89.47 | 24.41 | 79.17 | 27.93 | 74.8 | ||||||||||||||||||||||||||
| e5-mistral-7b-instruct | 56.89 | 61.88 | 38.35 | 42.97 | 48.89 | 87.84 | 56.59 | 75.72 | 43.06 | 38.62 | 63.53 | 89.61 | 16.3 | 76.41 | 26.39 | 87.25 | ||||||||||||||||||||||||||
| google-gecko | 55.7 | 62.18 | 33.21 | 48.89 | 47.12 | 86.96 | 59.24 | 71.33 | 32.58 | 40.33 | 61.28 | 88.18 | 20.34 | 75.42 | 25.86 | 82.62 | ||||||||||||||||||||||||||
| text-embedding-3-large | 55.44 | 58.05 | 30.27 | 47.54 | 44.76 | 87.94 | 55 | 71.58 | 40.24 | 42.07 | 61.27 | 89.05 | 23.11 | 77.77 | 23.35 | 79.56 | ||||||||||||||||||||||||||
| gte-en-large-v1.5 | 57.91 | 72.11 | 48.36 | 42.16 | 46.3 | 93.81 | 63.23 | 68.18 | 42.93 | 36.95 | 56.08 | 89.67 | 26.35 | 82.43 | 22.55 | 77.49 | ||||||||||||||||||||||||||
| gte-en-base-v1.5 | 54.09 | 63.49 | 40.36 | 39.52 | 39.9 | 94.81 | 48.65 | 67.75 | 42.62 | 35.88 | 52.96 | 88.42 | 21.92 | 76.77 | 25.22 | 73.13 | ||||||||||||||||||||||||||
| Multimodal Embedder | ||||||||||||||||||||||||||||||||||||||||||
| VISTA | 53.24 | 63.61 | 31.17 | 42.35 | 40.77 | 86.29 | 40.65 | 72.6 | 41.35 | 37.39 | 54.15 | 88.9 | 21.73 | 74.04 | 25.7 | 77.9 | ||||||||||||||||||||||||||
| CLIP-SF | 36.77 | 52.45 | 20 | 30.61 | 26.37 | 50.58 | 22.14 | 41.33 | 22.15 | 27.05 | 25.45 | 81.63 | 14.75 | 55.98 | 17.47 | 63.60 | ||||||||||||||||||||||||||
| One-Peace | 42.19 | 32.93 | 20.27 | 41.32 | 32.43 | 51.91 | 36.79 | 46.51 | 36.55 | 31.6 | 42.87 | 87.46 | 21.64 | 64.51 | 16.9 | 69.28 | ||||||||||||||||||||||||||
| DSE | 46.60 | 53.46 | 19.79 | 36.51 | 40.75 | 80.12 | 36.2 | 70.79 | 37.73 | 32.82 | 52.97 | 85.84 | 15.66 | 68.97 | 14.50 | 52.98 | ||||||||||||||||||||||||||
| E5-V | 49.91 | 54.28 | 21.64 | 41.69 | 38.78 | 78.99 | 45.41 | 60.88 | 41.23 | 36.97 | 51.58 | 87.6 | 22.36 | 72.75 | 21.61 | 72.85 | ||||||||||||||||||||||||||
| GME-Qwen2-VL-2B | 53.31 | 61.52 | 42.3 | 38.13 | 46.31 | 92.6 | 45.3 | 72.93 | 40.88 | 37.2 | 60.01 | 87.24 | 23.17 | 63.82 | 29.06 | 59.24 | ||||||||||||||||||||||||||
| GME-Qwen2-VL-7B | 55.68 | 64.60 | 45.38 | 41.66 | 50.78 | 94.27 | 57.14 | 79.21 | 42.38 | 38.40 | 67.74 | 88.05 | 27.38 | 62.31 | 23.26 | 52.6 | ||||||||||||||||||||||||||
8.3 Detailed Results on BEIR
BEIR is a heterogeneous benchmark containing diverse text IR tasks. We utilize BEIR to compare the performance of our GME with other text embedders on TT tasks. Table 9 presents the detailed evaluation nDCG@10 scores for pure text embedders and multimodal embedders on TT tasks. Except for our GME, other multimodal embedders do not match the performance of pure text embedders on text retrieval tasks, including those like E5-V that are fine-tuned exclusively on text data.
Naturally, pure text embedding models of the same model size still outperform multimodal embedding models in pure text retrieval tasks. For example, the score of the gte-Qwen2-7B-instruct model is 60.25, while the GME-Qwen2-VL-7B model, with the same model scale, scores 55.63. Although both models share the same text LLM, incorporating or extending multimodal capabilities leads to additional compromises in pure text performance. Minimizing this kind of loss remains an important research question.
| MBEIR | Avg. | qt | (,) | (,) | (,) | (,)(,) | ||||||||||||||||||||||||||||||||
EDIS |
NIGHTS | OVEN |
CIRR | OVEN |
| CLIP | 32.5 | 43.3 | 61.1 | 6.6 | 36.2 | 43.3 | 45.1 | 41.3 | 79.0 | 7.7 | 26.1 | 24.2 | 20.5 | 7.0 | 13.2 | 38.8 | 26.4 | |||||||||||||||||||||
| SigLIP | 37.2 | 30.1 | 75.7 | 36.5 | 39.8 | 27.0 | 43.5 | 30.8 | 88.2 | 34.2 | 28.9 | 29.7 | 25.1 | 14.4 | 22.7 | 41.7 | 27.4 | |||||||||||||||||||||
| BLIP | 26.8 | 16.4 | 74.4 | 15.9 | 44.9 | 26.8 | 20.3 | 17.2 | 83.2 | 19.9 | 27.4 | 16.1 | 10.2 | 2.3 | 10.6 | 27.4 | 16.6 | |||||||||||||||||||||
| BLIP2 | 24.8 | 16.7 | 63.8 | 14.0 | 38.6 | 26.9 | 24.5 | 15.0 | 80.0 | 14.2 | 25.4 | 12.2 | 5.5 | 4.4 | 11.8 | 27.3 | 15.8 | |||||||||||||||||||||
| VISTA | 26.37 | 5.77 | 47.97 | 3.08 | 83.80 | 36.78 | 54.84 | 2.79 | 48.92 | 4.72 | 24.43 | 22.32 | 18.53 | 3.28 | 14.65 | 27.77 | 22.27 | |||||||||||||||||||||
| CLIP-SF | 50.26 | 42.80 | 80.75 | 18.38 | 84.44 | 54.09 | 78.42 | 42.67 | 91.94 | 18.10 | 31.42 | 45.98 | 27.58 | 24.53 | 45.25 | 68.83 | 49.05 | |||||||||||||||||||||
| One-Peace | 38.00 | 48.95 | 71.45 | 32.34 | 63.67 | 53.01 | 32.42 | 47.27 | 85.60 | 30.89 | 31.27 | 23.69 | 20.05 | 2.93 | 10.53 | 30.56 | 23.32 | |||||||||||||||||||||
| DSE | 28.89 | 14.12 | 74.62 | 3.08 | 83.95 | 41.26 | 66.99 | 8.74 | 82.06 | 3.91 | 27.36 | 0.38 | 3.06 | 9.81 | 36.52 | 0.39 | 5.96 | |||||||||||||||||||||
| E5-V | 35.09 | 29.46 | 52.38 | 3.78 | 89.94 | 49.62 | 49.62 | 29.54 | 86.40 | 4.62 | 27.92 | 14.40 | 12.69 | 3.73 | 13.19 | 54.46 | 39.69 | |||||||||||||||||||||
| GME-Qwen2-VL-2B | 53.54 | 38.85 | 71.82 | 25.83 | 95.19 | 70.32 | 83.15 | 38.32 | 84.12 | 27.57 | 29.86 | 58.17 | 39.06 | 27.5 | 46.83 | 75.98 | 44.21 | |||||||||||||||||||||
| GME-Qwen2-VL-7B | 54.50 | 46.54 | 75.14 | 31.82 | 95.85 | 77.29 | 84.59 | 45.54 | 64.90 | 34.20 | 31.89 | 63.41 | 43.14 | 31.43 | 53.69 | 80.30 | 58.80 | |||||||||||||||||||||
8.4 Detailed Results on M-BEIR
M-BEIR, a multimodal benchmark for IR, serves as a comprehensive large-scale retrieval benchmark designed to evaluate multimodal retrieval models. As shown in Table 10, we report Recall@10 scores for the Fashion200K and FashionIQ datasets, while Recall@5 scores are provided for all other datasets. In M-BEIR, our GME continues to demonstrate state-of-the-art performance, underscoring the effectiveness of our approach.
| ArxivQ | DocQ | InfoQ | TabF | TATQ | Shift | AI | Energy | Gov. | Health. | Avg. | |
| BM25Text + Captioning | 40.1 | 38.4 | 70.0 | 35.4 | 61.5 | 60.9 | 88.0 | 84.7 | 82.7 | 89.2 | 65.1 |
| BGE-M3Text + Captioning | 35.7 | 32.9 | 71.9 | 69.1 | 43.8 | 73.1 | 88.8 | 83.3 | 80.4 | 91.3 | 67.0 |
| Jina-CLIP | 25.4 | 11.9 | 35.5 | 20.2 | 3.3 | 3.8 | 15.2 | 19.7 | 21.4 | 20.8 | 17.7 |
| Nomic-vision | 17.1 | 10.7 | 30.1 | 16.3 | 2.7 | 1.1 | 12.9 | 10.9 | 11.4 | 15.7 | 12.9 |
| SigLIP (Vanilla) | 43.2 | 30.3 | 64.1 | 58.1 | 26.2 | 18.7 | 62.5 | 65.7 | 66.1 | 79.1 | 51.4 |
| ColPali | 79.1 | 54.4 | 81.8 | 83.9 | 65.8 | 73.2 | 96.2 | 91.0 | 92.7 | 94.4 | 81.3 |
| VISTA | 10.3 | 8.01 | 30.02 | 13.08 | 2.05 | 3.26 | 7.14 | 11.05 | 6.9 | 9.39 | 10.12 |
| CLIP-SF | 24.1 | 11.8 | 48.78 | 21.53 | 5.49 | 6.06 | 28.64 | 27.19 | 34.67 | 32.64 | 24.09 |
| One-Peace | 43.94 | 23.48 | 59.97 | 57.05 | 13.44 | 17.02 | 45.41 | 53.21 | 55.98 | 59.5 | 42.9 |
| DSE | 78.17 | 45.83 | 82.06 | 79.29 | 49.01 | 69.84 | 96.89 | 92.62 | 92.04 | 96.35 | 78.21 |
| E5-V | 41.16 | 24.37 | 49.5 | 58.22 | 9.08 | 13.26 | 46.18 | 57.77 | 53.05 | 59.61 | 41.22 |
| GME-Qwen2-VL-2B | 83.91 | 54.57 | 91.11 | 94.61 | 71.05 | 94.29 | 99.02 | 93.15 | 97.89 | 98.89 | 87.84 |
| GME-Qwen2-VL-7B | 87.58 | 56.63 | 92.39 | 94.58 | 76.12 | 97.26 | 99.63 | 95.89 | 99.5 | 99.63 | 89.92 |
8.5 Detailed Results on ViDoRe
ViDoRe represents the Visual Document Retrieval Benchmark, encompassing various page-level screenshot retrieval tasks. This benchmark includes the TVD tasks within our UMRB. Table 11 presents the detailed nDCG@5 scores for our GME and other models. Our smaller model, i.e. GME-Qwen2-VL-2B, surpasses the previous state-of-the-art model (ColPali), which was exclusively trained on this dataset for this specific task. The larger model, i.e. GME-Qwen2-VL-7B, further improves upon this performance.
9 Experiment Details
| Hyper-param | GME-Qwen2-VL-2B | GME-Qwen2-VL-7B |
| Number of Params | 2B | 8.2B |
| Number of Layers | 28 | 28 |
| Hidden Size | 1536 | 3584 |
| FFN Inner Size | 3072 | |
| Number of Attention Heads | 12 | 28 |
| Vision Depth | 32 | |
| Vision Embed_dim | 1280 | |
| Vision Patch_size | 14 | |
| Temperature | 0.03 | |
| Learning Rate Decay | Linear | |
| Adam | 1e-4 | |
| Adam | 0.9 | |
| Adam | 0.98 | |
| Gradient Clipping | 0.0 | |
| Precision | PyTorch BF16 AMP | |
| Max Length | 1800 | 1800 |
| Batch Size | 128 | 32 |
| Warm-up Ratio | 0.06 | |
9.1 Training Details
Our GME models (both 2B and 7B) are initialized using the Qwen2-VL [65] model series. We employ the transformers library for training in BF16 precision. The training utilizes Low-Rank Adaptation (LoRA) [19] with a rank of 8. We apply a decoupled AdamW optimizer with a learning rate and a weight decay of 1e-4. Additional hyperparameters are detailed in Table 12.
In our contrastive learning approach, we develop dense multimodal representation models (embedders) that utilize the [EOS] hidden state as the embedding of the input. The temperature for contrastive learning is set to 0.03. For each query, we include one positive candidate along with eight hard negative candidates.
| Task | Dataset | Query Instruction |
| TT | ArguAna | Given a claim, find documents that refute the claim. |
| Climate-FEVER | Given a claim about climate change, retrieve documents that support orrefute the claim. | |
| CQADupStack | Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question. | |
| DBPedia | Given a query, retrieve relevant entity descriptions from DBPedia. | |
| FEVER | Given a claim, retrieve documents that support or refute the claim. | |
| FiQA2018 | Given a financial question, retrieve user replies that best answer the question. | |
| HotpotQA | Given a multi-hop question, retrieve documents that can help answer the question. | |
| MSMARCO | Given a web search query, retrieve relevant passages that answer the query. | |
| NFCorpus | Given a question, retrieve relevant documents that best answer the question. | |
| NQ | Given a question, retrieve Wikipedia passages that answer the question. | |
| Quora | Given a question, retrieve questions that are semantically equivalentto the given question. | |
| SCIDOCS | Given a scientific paper title, retrieve paper abstracts that are cited bythe given paper. | |
| SciFact | Given a scientific claim, retrieve documents that support or refute theclaim. | |
| Touche2020 | Given a question, retrieve detailed and persuasive arguments that answer the question. | |
| TRECCOVID | Given a query on COVID-19, retrieve documents that answer the query. | |
| WebQA | Retrieve passages from Wikipedia that provide answers to the following question. | |
| II | Nights | Find a day-to-day image that looks similar to the provided image. |
| TI | VisualNews | Identify the news-related image in line with the described event. |
| Fashion200k | Based on the following fashion description, retrieve the best matching image. | |
| MSCOCO | Identify the image showcasing the described everyday scene. | |
| Flickr30k | Find an image that matches the given caption. | |
| TVD | TAT-DQA | Find a screenshot that relevant to the user’s question. |
| ArxivQA | ||
| DocVQA | ||
| InfoVQA | ||
| Shift Project | ||
| Artificial Intelligence | ||
| Government Reports | ||
| Healthcare Industry | ||
| Energy | ||
| TabFQuad | ||
| IT | VisualNews | Find a caption for the news in the given photo. |
| Fashion200k | Find a product description for the fashion item in the image. | |
| MSCOCO | Find an image caption describing the following everyday image. | |
| Flickr30k | Find an image caption describing the following image. | |
| TIT | WebQA | Find a Wikipedia image that answers this question. |
| EDIS | Identify the news photo for the given caption. | |
| ITT | OVEN | Retrieve a Wikipedia paragraph that provides an answer to the given query about the image. |
| INFOSEEK | ||
| ReMuQ | Retrieve a fact-based paragraph that provides an answer to the given query about the image. | |
| OKVQA | Retrieve documents that provide an answer to the question alongside the image. | |
| LLaVA | Provide a specific decription of the image along with the following question. | |
| ITI | FashionIQ | Find a fashion image that aligns with the reference image and style note. |
| CIRR | Retrieve a day-to-day image that aligns with the modification instructions of the provided image. | |
| ITIT | OVEN | Retrieve a Wikipedia image-description pair that provides evidence for the question of this image. |
| INFOSEEK | ||
| EVQA | Obtain illustrated documents that correspond to the inquiry alongside the provided image. |
9.2 Instructions
The complete UMRB consists of 47 tasks, each with distinct retrieval candidates and varying domains. Even within the same dataset, retrieval candidates can differ based on task types. For example, the WebQA dataset aims to retrieve textual candidates for TT tasks, which is different from retrieving a combination of image and text candidates for TIT tasks.
We have designed specific instructions tailored for each task to guide the model in effectively completing the retrieval process. The detailed instructions are provided in Table 13.
10 Fused-Modal Data Synthesis Details
We utilize doc2query to synthesize data. However, our goal is to generate fused-modal candidate-to-query relevance data rather than single-modality, text-based relevance pairs.
10.1 Prompts
Step 1:
In the first step of data synthesis, we prompt the large language model (LLM) to generate a natural question and answer based on a selected passage. The specific prompt is illustrated in Figure 7. This process leverages in-context learning (ICL) to guide the LLM in producing outputs that align with our requirements.
Step 2:
In step 2, we provide the LLM with the passage and the natural question generated in step 1. The LLM is then prompted to extract the main entity from the question and refactor the question accordingly. Figure 8 presents the prompt used in this step. In subsequent steps, the extracted entity will be replaced by the corresponding image, which, when combined with the reconstructed question, will form a fused-modal query.
Step 3:
In step 3, we replace the entity with an image, which can be sourced in two ways. The first method involves prompting the LLM to generate a caption for the entity based on the provided entity and passage, after which the caption is fed into FLUX to generate images. The second method retrieves the entity by utilizing the Google Image Retrieval API. Figure 9 illustrates the caption generation prompt for this step.
10.2 Filter
Two filtering methods are implemented to ensure the quality of the synthesized data. First, a text retrieval model is utilized to evaluate unreconstructed queries and their corresponding passages. We follow the framework of Promptagator [9]; a query is deemed unqualified if the passage that generated it does not appear within the top 20 search results. Second, for images obtained through the Google Image Search API, we employ the CLIP model to assess image-caption relevance. Images with a relevance score below 0.2 are filtered out.
Why is the threshold score set to 0.2? The relevance scores of all images searched via Google and the corresponding captions we have collected are presented in Figure 10. We select the median score of 0.2 to ensure image quality while also ensuring that most text queries have sufficient images to pair with.
| Type | Task | Query Text | Query Image | Target Text | Target Image | |||||||||
| Single-Modal | TT |
| II | - |
| Cross-Modal | TI |
| TVD |
| IT |
| Fused-Modal | TIT |
| ITT |
| ITI |
| ITIT |
10.3 Examples of synthetic data
Table 15 illustrates passages from 15 domains and the fused modal queries generated by applying the synthesis flow. “FLUX image” refers to images generated by the Vincennes diagram model FLUX.1-dev, whereas “Google image” indicates images from Google Image retrieval.
| Domain | Candidate Image | Candidate Text | FLUX Image | Google Image | Query Text | |||||||||||
| animal |
| architecture |
| artwork |
| currency |
| entertainment |
| food |
| language |
| literature |
| mythology |
| organization |
| person |
| pharmaceutical |
| plant |
| sport |
| vehicle |
11 Limitations
In this work, we present a benchmark for training and testing Universal Multimodal Retrieval (UMR). To better accomplish this task, we explore strategies for adapting Multimodal Large Language Models (MLLMs) into UMR models, presenting GME, a powerful embedding model capable of retrieving candidates across different modalities. However, this work has its limitations, which are outlined below:
1. Single Image Limit
In MLLMs, one image is converted into a very large number of visual tokens. In Qwen2-VL, we limit the number of visual tokens to 1024. Due to model training efficiency and a lack of relevant data, our queries and candidates in UMRB only retain a single image. Thus, performance on interleaved data (where multiple images and texts are mixed together) cannot be assessed.
2. Single Language Limit
Although the backbone of our model, Qwen2-VL, supports multiple languages, we only utilized a single language, English, during the training and testing processes of our GME. Consequently, performance in other languages could not be evaluated.
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/night_q.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/night_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/fashion200k_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/vidore_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/visualnews_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/webqa_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/oven_q.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/fashioniq_q.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/fashioniq_c.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/evqa0.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/evqa_kb0.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/animal_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/animal_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/animal_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/architecture_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/architecture_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/architecture_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/artwork_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/artwork_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/artwork_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/currency_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/currency_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/currency_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/entertainment_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/entertainment_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/entertainment_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/food_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/food_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/food_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/language_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/language_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/language_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/literature_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/literature_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/literature_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/mythology_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/mythology_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/mythology_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/organization_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/organization_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/organization_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/person_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/person_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/person_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/pharmaceutical_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/pharmaceutical_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/pharmaceutical_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/plant_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/plant_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/plant_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/sport_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/sport_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/sport_google.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/vehicle_wiki.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/vehicle_flux.png)
![[Uncaptioned image]](extracted/6326670/figures/dataset_examples_figures/vehicle_google.png)